OpenStudy (anonymous):

the center of circle P is located at (5,-2) ST is a diameter of circle P. point T is located (-3,3). what are the coordinates of point S??

OpenStudy (ranga):

Since ST is a diameter, the center of the circle will be at the midpoint of S and T.

OpenStudy (ranga):

Let (a,b) be the coordinates of point S. T is (-3,3) The midpoint of ST is: ( (a-3)/2, (b+3)/2 ) which should be the center of the circle (5, -2) Equate the x-coordinates and solve for a: (a-3)/2 = 5; a-3 = 10; a = 13 Equate the y-coordinates and solve for b: (b+3)/2 = -2; b+3 = -4; b = -7 Coordinates of S is (13, -7).